# References.bib

### v1

@@ 4.1.4 Non-AI Threats and Defenses
Referring to our Figure 3, we specifically focus on DDoS attacks and supply chain issues. Even though these attacks are common problems among technology as a whole; they are threats we consider relevant to BC4LLMs. Ibrahim et al. [ibrahim_ddos_2022] suggest using a public blockchain to prevent DDoS attacks on IoT devices. Blockchain provides a tamper-proof platform as well as demonstrates how IoT devices working with blockchain can verify and authenticate using a trusted white-list which is implemented in the smart contract. Following this smart contract usage, if LLMs were to use a trusted white list for users then we can try to prevent these malicious users from trying to access the LLM in certain circumstances that are mutually agreed upon. Additionally proven in Shah et al. [shah_blockchain_2022], blockchain-based solutions play a vital role in mitigating DDoS attacks.
@@
A point of consideration is how DDoS attacks that target the blockchain to make the blockchain unavailable would require sufficient computer resources. The fully decentralized architecture of the blockchain and the consensus protocol for new blocks ensure that the blockchain can still operate meanwhile several blockchain nodes could be offline [zhang_securityandprivacyonblockchain_2020]. Incorporating this architecture into LLMs would help prevent DDoS attackers, as the larger the blockchain network is, then the harder it would be for a DDoS attack to be successful. Moreover, blockchain is known as a distributed, immutable, and verifiable ledger technology that ensures transparency and traceability [saberi_blockchain_2019]. By utilizing blockchain for LLMs, we can help mitigate these supply chain vulnerabilities. The decentralization of the network can maintain the integrity of the system at all points, aiding in mitigating the risk of a single point of failure, a common problem with centralized systems [saberi_blockchain_2019]. Blockchain is offered as a solution if the LLM were to accidentally crash, or was purposefully attacked by an attempt at overwhelming the system, then the LLM would still be intact since it is blockchain-based, removing the single point of failure entirely. However, it is important to note that blockchain solutions for LLMs depend on the availability of the underlying LLM infrastructure. If the LLM server is malfunctioning or shut down, then these blockchain mechanisms may not be applicable, highlighting the need for a robust and resilient supply chain. To solidify the supply chain, blockchain offers secure transactional data in sectors including supply chain management, healthcare, and federated learning [zuo_federatedtrustchain_2024]. For better supply chain management and data traceability, Kayikci and Khosgoftaar [kayikci_blockchain_2024] address the potential intersection of blockchain and ML. ML can aid in analyzing data from multiple sources and identify potential supply chain issues such as delays or quality issues before they occur. By using blockchain to create a transparent record of all supply chain transactions there are improvements in security, openness, traceability, and productivity [kayikci_blockchain_2024]. While blockchain presents a promising solution for enhancing security and defending against adversarial threats to LLMs, ongoing research and development are necessary to address the evolving landscape of threats and vulnerabilities.
4.2 Blockchain for Large Language Model Safety
@@
The burgeoning dominance of large language models as search engines [reid_generative_2024], code-writers [zheng_survey_2024], and a vast array of other roles has led to the emergence of unique problems within the sphere of their safety. For instance, LLMs who advise users to engage in dangerous activities such as eating glass [harrer_attention_2023] or which easily reveal personally identifying information [kim_propile_2023] may be unsafe for users to interact with even in the absence of external threats. In this section, we rely on our proposed definition of safety (Definition 2) to explore relevant literature that incorporates blockchain technology into the various solutions surrounding LLM safety.
4.2.1 Blockchain for Passive Privacy in Large Language Models
@@
Despite its novelty, the notion of passive privacy is of paramount importance in maintaining the safety of large language models. Some models face the issue of leaking sensitive data, potentially revealing private information like government-issued ID numbers and patient data  [pan_privacy_2020]. Given the severity of such leaks, it is evident that solutions to such problems are required for more substantial advancements in LLMs. To this end, blockchain’s guarantees about data sovereignty, obfuscation, and traceability translate well into realizable passive privacy benefits for large language models. In particular, we observe blockchain-based privacy preservation techniques which originate in varying proximity to LLMs as seen in BC4LLMs itself [wellington_basedai_2024][singh_enhancing_2024][ullah_privacy_2023], blockchain-enabled deep learning [shafay_blockchain_2023][zhu_blockchain-based_2019][weng_deepchain_2021][weng_auditable_2023][keshk_privacy-preserving-framework-based_2020], blockchain-enabled machine learning [arachchige_trustworthy_2020], and blockchain-enabled federated learning [qammar_securing_2023][m_privacy-preserving_2024][zhao_privacy-preserving_2021][nagar_privacy-preserving_2019][awan_poster_2019][ferenczi_fully_2023][qu_blockchain-enabled_2023].
@@
Within our focus of BC4LLMs, we have observed distinct trends in the application of blockchain to large language models in their capacity to bolster passive privacy guarantees. Most notably, the development of ZK-LLMs, or zero-knowledge large language models as described in Wellington [wellington_basedai_2024] and Singh [singh_enhancing_2024] have the potential to drastically reduce privacy leakage risks when interacting with large language models. Considering the problem of data leakage approached from the lens of access, this application is natural. A user querying for their own personally identifiable information should, ideally, be able to access it whereas an unauthorized user should not. Obfuscating portions or the entirety of a corpus using zero-knowledge proofs allows for untrusted training nodes, or the model itself, to act on sensitive data without the ability to regurgitate it to a potentially malicious party. This same mechanism has broad applications that have been explored in other recent works as well, with special focuses on ZKPs for data curation and pre-processing [ullah_privacy_2023] which consequently enhance passive privacy within large language models.
@@
Additionally, besides material on BC4LLMs, it is necessary to discuss passive privacy contributions made within LLM-related areas, as described in our classification of LLMs in the context of AI, ML, and DL.(Figure 2) Especially important in its immediate applications to LLMs, blockchain-enabled deep learning (BC-DL) is a growing field with potentially large impacts on LLM’s passive privacy. Specifically, certain BC-DL technologies propose learning mechanisms distinct from traditional federated learning models [weng_deepchain_2021][shafay_blockchain_2023][keshk_privacy-preserving-framework-based_2020]. The concerted research effort to develop efficient distributed learning models that deviate from the typical model of federated learning is clearly well underway. This field has broad implications for blockchain; through the utilization of various blockchain properties, we see the development of privacy guarantees which undoubtedly strengthens the BC4LLMs area.
Of particular note in the BC-DL area is the influential DeepChain paper [weng_deepchain_2021] which introduced a novel privacy-preserving training framework based on blockchain technology. The proposed system, reliant on a consensus protocol and corresponding incentive mechanism, allows for the existence of private training gradients and the guaranteed auditability of training data. This type of twofold approach, which also leans on zero knowledge in several areas of the protocol, could be indicative of a new direction of research for blockchain-enabled passive privacy, stemming from the vastly studied federated learning area. In order to highlight the novelty of such an approach, we provide Figure 6 to further illuminate the nature of a system interacted with chiefly through the two vehicles of consensus and incentivization.
@@
Figure 6: Within DeepChain, we see a deviation from conventional federated learning models in that while training occurs in iterative rounds, it is marked by synchronous requirements for consensus on data added to the model in the last round of training. The contribution-based incentive mechanism rewards workers for verifying new gradients, and stimulates the trading contract which facilitates the sharing of new and updated gradient information. A Verifiable Random Function (VRF) ensures that the committee-based consensus is fair and alleviates concerns surrounding finality.
@@
Additionally, despite this potential research direction, discussion on the wide body of research that does exist concerning blockchain-enabled federated learning (BC-FL) is essential when describing blockchain as a vehicle for improving the safety and security of large language models. Federated learning, introduced in McMahan et al. [mcmahan_communication-efficient_2017], has since been the principle building block of decentralized learning approaches in machine learning systems. While too broad to be considered for BC4LLMs, notable to this paper are the contributions of authors approaching BC-FL in its capacity as a powerful privacy preserving mechanism [nagar_privacy-preserving_2019][awan_poster_2019][ferenczi_fully_2023][fotohi_decentralized_2024][m_privacy-preserving_2024][zhao_privacy-preserving_2021].
4.2.2 Blockchain for Misinformation in Large Language Models
@@
The issue of large language models fabricating information, commonly known as hallucinations [maynez_faithfulness_2020][openai_gpt-4_2024] is well understood. As a result, detecting and defending against hallucinations is a widely explored area [seneviratne_blockchain_2022], with more research still yet to be conducted [ye_cognitive_2023][zhang_sirens_2023]. Along with this pressing general body of research, efforts have been made to leverage blockchain technology to reduce hallucinations by consensus-oriented [zhang_proof_2024] and oracle-based [bouchiha_llmchain_2024] approaches. Within consensus, Zhang et al. [zhang_proof_2024] proposed a system for efficient large language model inference quality assessment. That is, the veracity of a given model’s responses was able to be assessed by using a ’Proof of Quality’ consensus mechanism with low latency between the user and language model. This stands in contrast to other approaches, such as Bouchiha et al.’s [bouchiha_llmchain_2024] reputation-based system LLMChain, which relies on a decentralized oracle to cross reference request/response pairs originating from differing models and speak to the quality of inferences based on those comparisons. It is worth noting that despite these fundamental differences, they are both consensus-based approaches. This serves as an excellent example of how BC4LLMs technology can take many different forms towards the same goal.
@@
Besides steering towards consensus-driven models of governance for misinformation, several other ideas have been considered to enhance the correctness of large language model responses. These accomplishments have come from the zero-knowledge space as in the case of Chen et al. [chen_zkml_2024], more optimistic privacy guarantees found within works such as Conway et al. [conway_opml_2024], and even straightforward applications of the verifiable ledger as in Yazdinejad et al. [yazdinejad_making_2020]. Chen et al. propose zkML, a compiler that enables TensorFlow models to be translated readily into zk-SNARK halo2 circuits via either KZG or IPA commitments. This conversion allows for any portion, or the entirety of, an LLM to gain the properties of zero-knowledge, knowledge soundness, and completeness. Through this, and with potential connections to verifiable databases, zkML gains the powerful ability to audit inferences and ensure their accuracy. This research avenue is particularly promising due to both the efficient and potentially on-chain verification of zk-SNARKs as well as the extensibility of zkML to virtually any ML model.
@@
Distinct from the zkML approach is that of Conway et al. with opML [conway_opml_2024], which opts for an optimistic approach reliant on a fraud proof rather than a ZK proof to catch erroneous outputs within a certain challenge period. Clearly, there exist trade-offs in this implementation when compared to the zkML approach. Optimistic rollups are desirable in the sense that they are performant, but if implemented in a RAG environment, or similarly situated between the user and a model, latency issues can quickly become dominant. Apart from proof-oriented mechanisms and worth noting is the work of Yazdinejad et al. [yazdinejad_making_2020], which focuses on detecting deep fakes using blockchain’s verifiable ledger. While not directly applicable to the realm of LLMs due to the non-atomic nature of data within a language model, important insights can be drawn from the paper. Namely, BC4LLMs could benefit greatly from a proposed hashing method applied to particularly sensitive data areas such as names, addresses, or even health-care related parts of corpora. This hash could be used as a guarantee of data veracity and could potentially prevent unsafe behaviors such as sycophancy, deception, or unfairness. Indeed, this hashing mechanism has the potential to be used as a final check for the LLM to verify that it is submitting information to the user that is consistent with standards agreed upon when information was originally committed to the ledger. Many similar vehicles for the maintenance of data integrity exist, albeit currently limited by scaling issues on-chain [deshpande_distributed_nodate].
@@ 5 Datasets Relevant to BC4LLM
Developing synergistic technologies that incorporate blockchain and large language models is vital to securing a future consisting of safe, trustworthy LLMs. To this end, relevant datasets with which researchers may conduct experiments are of paramount importance. Moreover, it is often the case that blockchain-enabled systems require unorthodox training sets and edge cases in order to fully capture the dynamism and robustness of respective implementations. Therefore we have collected and summarize the relevance of certain datasets in Table 4. Of note are the standard datasets such as MNIST, CIFAR-10, SQuAD, and MS-MARCO, and we also give special attention to less widely known datasets which still may prove useful in certain academic settings.
@@
MNIST11footnotemark: 1 Pattern Recognition Images of handwritten digits for pattern recognition applications oe vulnerability analysis. [weng_deepchain_2021][weng_mystique_2021][mcmahan_communication-efficient_2017][ferenczi_fully_2023][shen_exploiting_2021] [li_byzantine_2021]
CIFAR-1022footnotemark: 2 Labeled images used in capacities from improving pattern recognition to zk-SNARK benchmarks. [weng_mystique_2021][chen_zkml_2024][mcmahan_communication-efficient_2017]
MS MARCO33footnotemark: 3 Collection of human answered questions, used in training corpora as well as simulating RAG attacks. [xue_badrag_2024][zou_poisonedrag_2024][cheng_trojanrag_2024]
MedMINST44footnotemark: 4 Collection of medical images from case studies. [xue_badrag_2024][zou_poisonedrag_2024][cheng_trojanrag_2024]
Natural Questions55footnotemark: 5 Poisoned RAG Open domain question answering dataset, incorporating questions from users and rigorous answers. [xue_badrag_2024][zou_poisonedrag_2024][cheng_trojanrag_2024]
HotpotQA66footnotemark: 6 Question answering dataset with multi-hop questions and supervised, regulated, answers. [zou_poisonedrag_2024][cheng_trojanrag_2024]
MT BENCH77footnotemark: 7 LLM Evaluation Ranked pairwise expert human preferences for various model responses. [zheng_judging_2023][bouchiha_llmchain_2024]
SQuAD88footnotemark: 8 Reading comprehension dataset comprised of questions posed on Wikipedia article with answers as sections of those corresponding articles. [zuo_federatedtrustchain_2024][hu_prompt_2024]
IMDB Dataset99footnotemark: 9 Sentiment Analysis Movie reviews [zuo_federatedtrustchain_2024][hu_prompt_2024]
SafetyBench1010footnotemark: 10 Safety Evaluation Large number of multiple choice questions focused on evaluating the safety of large language models. [zhang_safetybench_2023]
@@
6 Prominent Challenges Within
BC4LLM
Despite the promise of the emerging BC4LLMs field, there are several innate challenges that delay progress and inhibit potential research directions. Typically, these are derived from certain limitations in blockchain technology, large language models, or deficits in the way that blockchain can serve large language models.

### v2

@@ 4.1.4 Non-AI Threats and Defenses
Referring to Figure 3, we specifically focus on DDoS attacks and supply chain issues. Even though these attacks are common problems, we consider threats relevant to BC4LLMs. Ibrahim et al. [ibrahim_ddos_2022] suggest using a public blockchain to prevent DDoS attacks on IoT devices. Blockchain provides a tamper-proof platform as well as demonstrates how IoT devices working with blockchain can verify and authenticate using a trusted white-list which is implemented in the smart contract. Following this smart contract usage, if LLMs were to use a trusted white list for users then we can try to prevent these malicious users from trying to access the LLM in certain circumstances that are mutually agreed upon. Additionally proven by Shah et al. [shah_blockchain_2022], blockchain-based solutions play a vital role in mitigating DDoS attacks.
@@
A point of consideration is how DDoS attacks that target the blockchain to make the blockchain unavailable would require sufficient computer resources. The fully decentralized architecture of the blockchain and the consensus protocol for new blocks ensure that the blockchain can still operate meanwhile several blockchain nodes could be offline [zhang_securityandprivacyonblockchain_2020]. Incorporating this architecture into LLMs would help prevent DDoS attackers, as the larger the blockchain network is, then the harder it would be for a DDoS attack to be successful. Moreover, blockchain is known as a distributed, immutable, and verifiable ledger technology that ensures transparency and traceability [saberi_blockchain_2019]. By utilizing blockchain for LLMs, we can help mitigate these supply chain vulnerabilities. The decentralization of the network can maintain the integrity of the system at all points, aiding in mitigating the risk of a single point of failure, a common problem with centralized systems [saberi_blockchain_2019]. Blockchain is offered as a solution if the LLM were to accidentally crash, or was purposefully attacked by an attempt at overwhelming the system, then the LLM would still be intact since it is blockchain-based, removing the single point of failure entirely. However, it is important to note that blockchain solutions for LLMs depend on the availability of the underlying LLM infrastructure. If the LLM server is malfunctioning or shuts down, then these blockchain mechanisms may not be applicable, highlighting the need for a robust and resilient supply chain. To solidify the supply chain, blockchain offers secure transactional data in sectors including supply chain management, healthcare, and federated learning [zuo_federatedtrustchain_2024]. For better supply chain management and data traceability, Kayikci and Khosgoftaar [kayikci_blockchain_2024] address the potential intersection of blockchain and ML. ML can aid in analyzing data from multiple sources and identify potential supply chain issues such as delays or quality issues before they occur. By using blockchain to create a transparent record of all supply chain transactions there are improvements in security, openness, traceability, and productivity [kayikci_blockchain_2024]. While blockchain presents a promising solution for enhancing security and defending against adversarial threats to LLMs, ongoing research and development are necessary to address the evolving landscape of threats and vulnerabilities.
4.2 Blockchain for LLMs’ Safety
@@
The growing dominance of LLMs as search engines [reid_generative_2024], code writers [zheng_survey_2023], and in many other roles has introduced unique challenges related to their safety. For instance, LLMs who advise users to engage in dangerous activities such as eating glass [harrer_attention_2023] or which easily reveal personally identifying information [kim_propile_2024] may be unsafe for users to interact with even in the absence of external threats. In this section, we rely on our proposed definition of safety (Definition 2) to explore relevant literature that incorporates blockchain technology into the various solutions surrounding LLM safety.
4.2.1 Blockchain for Passive Privacy in LLMs
@@
Despite its novelty, the concept of passive privacy is crucial for ensuring the safety of LLMs. Some models risk leaking sensitive information, potentially exposing private data like government-issued ID numbers and patient records [pan_privacy_2020]. The severity of these leaks underscores the need for effective solutions to advance LLMs responsibly. In this regard, blockchain’s guarantees of data sovereignty, obfuscation, and traceability offer practical passive privacy benefits that align well with the requirements of LLMs. In particular, we observe blockchain-based privacy preservation techniques which originate in varying proximity to LLMs as seen in BC4LLMs itself [ullah_privacy_2023, wellington_basedai_2024, singh_enhancing_2024], blockchain-enabled deep learning [zhu_blockchain-based_2019, keshk_privacy-preserving-framework-based_2020, weng_deepchain_2021, weng_auditable_2023, shafay_blockchain_2023], blockchain-enabled machine learning [arachchige_trustworthy_2020], and blockchain-enabled federated learning [awan_poster_2019, nagar_privacy-preserving_2019, zhao_privacy-preserving_2021, qammar_securing_2023, ferenczi_fully_2023, qu_blockchain-enabled_2023, m_privacy-preserving_2024].
@@
Within our focus on BC4LLMs, we have observed distinct trends in the application of blockchain to LLMs in their capacity to bolster passive privacy guarantees. Most notably, the development of zero-knowledge LLMs, a.k.a. ZK-LLMs, as described in [wellington_basedai_2024] and [singh_enhancing_2024], has the potential to drastically reduce privacy leakage risks when interacting with LLMs. Considering the problem of data leakage approached from the lens of access, this application is natural. A user querying for their own personally identifiable information should, ideally, be able to access it whereas an unauthorized user should not. Obfuscating portions or the entirety of a corpus using zero-knowledge proofs [zhang2020transparent] allows for untrusted training nodes, or the model itself, to act on sensitive data without the ability to regurgitate it to a potentially malicious party. This same mechanism has broad applications that have been explored in other recent works as well, with special focuses on ZKPs for data curation and pre-processing [ullah_privacy_2023], which consequently enhance passive privacy within LLMs.
@@
Additionally, besides material on BC4LLMs, it is necessary to discuss passive privacy contributions made within LLM-related areas, as described in our classification of LLMs in the context of AI, ML, and DL (Figure 2). Especially important in its immediate applications to LLMs, blockchain-enabled deep learning (BC-DL) is a growing field with potentially large impacts on LLM’s passive privacy. Specifically, certain BC-DL technologies propose learning mechanisms distinct from traditional federated learning models [keshk_privacy-preserving-framework-based_2020, weng_deepchain_2021, shafay_blockchain_2023]. The concerted research effort to develop efficient distributed learning models that deviate from the typical model of federated learning is clearly well underway. This field has broad implications for blockchain; through the utilization of various blockchain properties, we see the development of privacy guarantees which undoubtedly strengthens the BC4LLMs area.
A noteworthy contribution in the field of blockchain and deep learning is the influential DeepChain [weng_deepchain_2021], which introduces a novel privacy-preserving training framework based on blockchain technology. This system employs a consensus protocol alongside an incentive mechanism, enabling the use of private training gradients and ensuring the auditability of training data. This dual approach, incorporating zero-knowledge proofs in various aspects of the protocol, represents a promising new direction for blockchain-enabled passive privacy, building upon the well-established domain of federated learning. To underscore the novelty of this approach, Figure 6 illustrates how the system operates primarily through consensus and incentivization mechanisms.
@@
Figure 6: DeepChain deviates from conventional federated learning models by incorporating a synchronous requirement for consensus on data incorporated into the model during the final training round. The contribution-based incentive mechanism rewards participants for verifying new gradients and activates a trading contract to facilitate the sharing of updated gradient information. Besides, a Verifiable Random Function (VRF) ensures fairness in the committee-based consensus process, addressing concerns related to finality.
@@
Additionally, despite this potential research direction, discussion on the wide body of research that does exist concerning blockchain-enabled federated learning (BC-FL) is essential when describing blockchain as a vehicle for improving the safety and security of LLMs. Federated learning, introduced by McMahan et al. [mcmahan_communication-efficient_2017], has since been the principal building block of decentralized learning approaches in machine learning systems. While too broad to be considered for BC4LLMs, notable to this paper are the contributions of authors approaching BC-FL in its capacity as a powerful privacy preserving mechanism [nagar_privacy-preserving_2019, awan_poster_2019, zhao_privacy-preserving_2021, ferenczi_fully_2023, fotohi_decentralized_2024, m_privacy-preserving_2024].
4.2.2 Blockchain for Misinformation in LLMs
@@
The issue of LLMs fabricating information, commonly known as hallucinations [maynez_faithfulness_2020, openai_gpt-4_2024] is well understood. As a result, detecting and defending against hallucinations is a widely explored area [seneviratne_blockchain_2022], with more research still yet to be conducted [ye_cognitive_2023, zhang_sirens_2023]. Along with this pressing general body of research, efforts have been made to leverage blockchain technology to reduce hallucinations by consensus-oriented [zhang_proof_2024] and oracle-based [bouchiha_llmchain_2024] approaches. Within consensus, Zhang et al. [zhang_proof_2024] proposed a system for efficient large language model inference quality assessment. That is, the veracity of a given model’s responses was able to be assessed by using a ‘Proof of Quality’ consensus mechanism with low latency between the user and language model. This stands in contrast to other approaches, such as Bouchiha et al.’s [bouchiha_llmchain_2024] reputation-based system LLMChain, which relies on a decentralized oracle to cross reference request/response pairs originating from differing models and speak to the quality of inferences based on those comparisons. It is worth noting that despite these fundamental differences, they are both consensus-based approaches. This serves as an excellent example of how BC4LLM technology can take many different forms towards the same goal.
@@
In addition to advancing consensus-driven governance models for mitigating misinformation, several approaches have been explored to improve the accuracy of LLMs’ responses. These efforts include contributions from the zero-knowledge domain, as demonstrated by Chen et al. [chen_zkml_2024], more optimistic privacy assurances found in works like [conway_opml_2024], and straightforward applications of verifiable ledgers, as proposed by Yazdinejad et al. [yazdinejad_making_2020]. Chen et al. [chen_zkml_2024] propose zkML, a compiler that enables TensorFlow models to be translated readily into zk-SNARK halo2 circuits via either KZG or IPA commitments. This conversion allows for any portion, or the entirety of, an LLM to gain the properties of zero-knowledge, knowledge soundness, and completeness. Through this, and with potential connections to verifiable databases, zkML gains the powerful ability to audit inferences and ensure their accuracy. This research avenue is particularly promising due to both the efficient and potentially on-chain verification of zk-SNARKs as well as the extensibility of zkML to virtually any ML model.
@@
Distinct from the zkML approach is opML [conway_opml_2024], which opts for an optimistic approach reliant on a fraud-proof rather than a ZK proof to catch erroneous outputs within a certain challenge period. Clearly, there exist trade-offs in this implementation when compared to the zkML approach. Optimistic rollups are desirable in the sense that they are performant, but if implemented in a RAG environment, or similarly situated between the user and a model, latency issues can quickly become dominant. Apart from proof-oriented mechanisms and worth noting is the work proposed by Yazdinejad et al. [yazdinejad_making_2020], which focuses on detecting deepfakes using blockchain’s verifiable ledger. While not directly applicable to the realm of LLMs due to the non-atomic nature of data within a language model, important insights can be drawn from the paper. Namely, BC4LLMs could benefit greatly from a proposed hashing method applied to particularly sensitive data areas such as names, addresses, or even health-care-related parts of corpora. This hash could be used as a guarantee of data veracity and could potentially prevent unsafe behaviors such as sycophancy, deception, or unfairness. Indeed, this hashing mechanism has the potential to be used as a final check for the LLM to verify that it is submitting information to the user that is consistent with standards agreed upon when information was originally committed to the ledger. Many similar vehicles for the maintenance of data integrity exist, albeit currently limited by scaling issues on-chain [deshpande_distributed_nodate].
@@ 5 Datasets Relevant to BC4LLM
Developing synergistic techniques that integrate blockchain with LLMs is essential for ensuring the safety and trustworthiness of future LLMs. In this context, it is critical to access to relevant datasets for experimentation. Blockchain-enabled systems often require unconventional training sets and edge cases to capture the dynamism and robustness of these implementations. Accordingly, we have compiled and summarized the relevance of specific datasets in Table 4. While we include standard datasets such as MNIST, CIFAR-10, SQuAD, and MS-MARCO, we also highlight lesser-known datasets that may prove valuable in particular research contexts.
@@
MNIST11footnotemark: 1 Pattern Recognition Images of handwritten digits for pattern recognition applications or vulnerability analysis. [weng_deepchain_2021, weng_mystique_2021, mcmahan_communication-efficient_2017, ferenczi_fully_2023, shen_exploiting_2021, li_byzantine_2021]
CIFAR-1022footnotemark: 2 Labeled images used in capacities from improving pattern recognition to zk-SNARK benchmarks. [weng_mystique_2021, chen_zkml_2024, mcmahan_communication-efficient_2017]
MS MARCO33footnotemark: 3 Collection of human answered questions, used in training corpora as well as simulating RAG attacks. [xue_badrag_2024, zou_poisonedrag_2024, cheng_trojanrag_2024]
MedMINST44footnotemark: 4 Collection of medical images from case studies. [xue_badrag_2024, zou_poisonedrag_2024, cheng_trojanrag_2024]
Natural Questions55footnotemark: 5 Poisoned RAG Open domain question answering dataset, incorporating questions from users and rigorous answers. [xue_badrag_2024, zou_poisonedrag_2024, cheng_trojanrag_2024]
HotpotQA66footnotemark: 6 Question answering dataset with multi-hop questions and supervised, regulated, answers. [zou_poisonedrag_2024, cheng_trojanrag_2024]
MT BENCH77footnotemark: 7 LLM Evaluation Ranked pairwise expert human preferences for various model responses. [zheng_judging_2023, bouchiha_llmchain_2024]
SQuAD88footnotemark: 8 Reading comprehension dataset comprised of questions posed on Wikipedia article with answers as sections of those corresponding articles. [zuo_federatedtrustchain_2024, hu_prompt_2024]
IMDB Dataset99footnotemark: 9 Sentiment Analysis Movie reviews [zuo_federatedtrustchain_2024, hu_prompt_2024]
SafetyBench1010footnotemark: 10 Safety Evaluation Large number of multiple choice questions focused on evaluating the safety of LLMs. [zhang_safetybench_2023]
@@
6 Challenges in BC4LLM
Despite the promise of the emerging BC4LLMs field, there are several innate challenges that delay progress and inhibit potential research directions. Typically, these are derived from certain limitations in blockchain technology, LLMs, or deficits in the way that blockchain can serve LLMs.